OpenStudy (anonymous):

2.Max was solving the equation below and isn’t sure if his answer is correct. Explain to Max how he can check his answer and then help him identify any errors he made. Provide the correct solution in your explanation. 3x+6 8 = 7x−1 6 24 1 ( 3x+6 8 ) = 24 1 ( 7x−1 6 ) 3(3x+6) = 4(7x−1) 9x+18 = 28x−1 9x−9x+18 = 28x−9x−1 18 = 19x−1 18+1 = 19x−1+1 19 = 19x 1 = x

OpenStudy (welshfella):

if you plug x = 1 into the original equation you will see that it doesn't satisfy LHS = 27 and RHS = 24

OpenStudy (anonymous):

im confused

OpenStudy (welshfella):

replace the x's by 1 3(3(1) + 6) = 3 * 9 = 27 4(7(1) - 1) = 4*6 = 24 left side does not equal right side so x = m1 is not a solution

OpenStudy (welshfella):

*x = 1

OpenStudy (anonymous):

okay